E10-5 Calculating Return on Investment, Residual Income, Determining Effect of Changes in Sales, Expenses, Invested Assets, Hd. Average invested assets increase by $500,000. Return on Investment Residual Income (Loss)
Answer #1

1) Calculate following

Return on investment (284000/2600000) 10.92%
Investment turnover (900000/2600000) 0.3462
Profit margin (284000/900000) 31.56%
Residual income (284000-312000) -28000

2) Return on investment = 284000/3100000 = 9.16%

Residual income = 284000-372000 = -88000